OpenStudy (followthefez):

On the basis of simple electronegativity, which atom in the following molecule would you expect to be most strongly delta-negative? CH3 - CH = CH - CHOH - CHCl - CH2 - NH2

OpenStudy (followthefez):

Actual molecule is drawn as follows. 

OpenStudy (abb0t):

-∆? ∆H? ∆G? ∆V? ∆S?

OpenStudy (followthefez):

\[\delta- \] <-that's what I meant by 'delta negative'

OpenStudy (anonymous):

do you still need help or have you gotten it?

OpenStudy (followthefez):

Need help. Actually I think I know what the answer is but the book says I'm wrong and my lecturer says that some of the answers in the book are incorrect. Just want to check if this is one of those answers. lol. If you can let me know what you think it is please so I can check??

OpenStudy (anonymous):

sure whats your books answer

OpenStudy (followthefez):

Book says it's O. But I'm pretty sure it's Cl.

OpenStudy (anonymous):

I think it could be O but Cl seems correct

OpenStudy (followthefez):

But you're not too sure which one if either is the correct answer??

OpenStudy (anonymous):

I think you're correct but go talk to a peer or something

OpenStudy (anonymous):

i'm guessing your a whovian :P

OpenStudy (aaronq):

i think it would be the oxygen, in terms of comparison between the O and H there is a greater electronegativity difference than between Cl and C. Cl - C = 3.16 - 2.55 < 3.44 -2.2 = O - H